# ExioML: Eco-economic dataset for Machine Learning in Global Sectoral Sustainability

### v1

@@ Abstract
The Environmental Extended Multi-Regional Input-Output analysis is the predominant framework in Ecological Economics for assessing the environmental impact of economic activities. This paper introduces ExioML, the first Machine Learning benchmark dataset designed for sustainability analysis, aimed at lowering barriers and fostering collaboration between Machine Learning and Ecological Economics research. A crucial greenhouse gas emission regression task was conducted to evaluate sectoral sustainability and demonstrate the usability of the dataset. We compared the performance of traditional shallow models with deep learning models, utilizing a diverse Factor Accounting table and incorporating various categorical and numerical features. Our findings reveal that ExioML, with its high usability, enables deep and ensemble models to achieve low mean square errors, establishing a baseline for future Machine Learning research. Through ExioML, we aim to build a foundational dataset supporting various Machine Learning applications and promote climate actions and sustainable investment decisions. The codes are available: https://github.com/Yvnminc/ExioML
1 Introduction
@@
The increase in Greenhouse Gas (GHG) emissions due to fossil-fuel-driven economic development has precipitated a global warming crisis. This concern led to establishing the Paris Agreement in 2015, aiming to limit long-term temperature rise to no more than 2 C above pre-industrial levels (Schleussner et al., 2016). Concurrently, the Sustainable Development Goals (SDGs), specifically Goal 13 and Goal 8, were proposed to emphasize taking climate action to reduce emissions while maintaining economic growth. To address such climate-trade dilemma, researchers from various disciplines strive to balance climate action with economic growth and human well-being (Rao et al., 2014; Jorgenson, 2014). Recently, the Machine Learning (ML) technique (LeCun et al., 2015) has emerged as a significant tool for accurate prediction to assist climate change decision-making (Rolnick et al., 2022). Specifically, ML algorithms have been explored in aiding nearly real-time global weather forecasting (Lam et al., 2023), land monitoring via satellite imagery (Stanimirova et al., 2023; He et al., 2016), and the prediction of disturbances in electric grids (Zheng et al., 2022).
However, unlike the open-source culture of the ML community (Zheng et al., 2022), most Ecological Economics (EE) studies haven’t publicised the code and data. This leads to fragmented, independent, and inconsistent research leveraging ML techniques to explore the sustainability of different sectors and regions. Subsequent research cannot fairly compare the model performance, such as accuracy, robustness and generalisation or reproduce the result based on previous research, leading to the slow EE-ML development (Zhu et al., 2023; Ballarin et al., 2023; Nangini et al., 2019). We proposed ExioML as the first ML-ready benchmark data in EE research. The ExioML is developed on top of the high-quality open-source EE-MRIO dataset ExioBase 3.8.2 with high spatiotemporal resolution (Stadler et al., 2018). Enabling tracking resource flow between international sectors for global sectoral sustainability assessment. The usability of the ExioML dataset was validated using multimodality regression to quantify the sectoral GHG emission, achieving a low MSE result. This establishes a critical baseline for accuracy and efficiency in future ML research.
2 Literature Review
@@
No public benchmark Ecological Economics and Machine Learning (EE-ML) dataset is available to the best of our knowledge. The EE-MRIO data is a high-quality source for constructing ML-ready datasets. However, the existing dataset faces limitations in terms of accessibility and resolution. Specifically, Eora offers high-quality data but is closed-source and requires high-cost purchasing (Lenzen et al., 2013). Global Trade Analysis Project (GTAP) covers only five reference time steps with non-free access (Chepeliev, 2023). The World Input-Output Database (WIOD) (Dietzenbacher et al., 2013) has been questioned for its low resolution of temporal and spatial scope, whose latest temporal coverage is 2016. Therefore, uniform problem formulations and benchmark datasets need to be urgently developed. In summary, there exists a big gap in the EE-ML cooperation, which is demonstrated in three aspects: the inaccessibility of the dataset, the intensive data preprocessing with domain knowledge are required and the lack of Benchmark ML datasets and models.
@@
The predominant Ecological Economics research framework, the Environmentally Extended Multiregional Input-Output (EE-MRIO) analysis, effectively models global economic interactions of sectors within a network structure (Leontief & Strout, 1963; Wang et al., 2019; Sun et al., 2020a; Steinberger et al., 2012; Jakob & Marschinski, 2013). As EE-MRIO describes the environmental footprint for global economic activities, it has become the fundamental framework for EE research, illustrated in Figure 1. EE-MRIO supports various studies such as Structure Decomposition Analysis (SDA) and Index Decomposition Analysis (IDA) to identify changes by decomposing key drivers (Hoekstra & Van den Bergh, 2003; Peters et al., 2017; Duan & Yan, 2019); monitoring embodied emissions with resource transfer and sustainability evaluation of supply chain in global trade (Meng et al., 2023; Tian et al., 2022; Wu et al., 2020; He & Hertwich, 2019; Li et al., 2020; Long et al., 2018; Sun et al., 2020a). Recently, ML algorithms have been applied with EE-MRIO for several applications, such as accurately identifying ecological hotspots and inefficiencies within the global supply chain to optimise logistic paths for decreased carbon emissions while considering cost-effectiveness (Akbari & Do, 2021). ML algorithms are naturally suitable for learning multi-dimensional patterns and are utilised for better sectoral sustainability assessment considering environmental, economic and social impacts (Abdella et al., 2020; Nilashi et al., 2019). Additionally, the ML algorithms extract the patterns from the original high-dimensional space to latent feature space, and it is leveraged to extract the critical paths of products’ energy, water, and emission footprints from giant trading networks for supporting sustainable investment decisions (Ding et al., 2022).
@@ 3 Methods
The architecture of the proposed ExioML dataset is illustrated in Figure 2. It ends the challenges of data limitation and simplifies the EE-MRIO framework for ML researchers in a ready-to-use manner. ExioML offers high spatiotemporal resolution for 49 regions from 1995 to 2022 containing tabular and graphical formats, supporting diverse ML research, such as time-series emissions forecasting (Deb et al., 2017), factor analysis (Raihan et al., 2022; Zhang et al., 2019; Duan & Yan, 2019; Acheampong & Boateng, 2019; Khan et al., 2020), graph learning (Sun et al., 2020b) and clustering analysis (He et al., 2022; Kijewska & Bluszcz, 2016), under a uniform framework. The details of ExioML are summarised in Tablue 1.
@@
ExioML is derived from an open-source EE-MRIO dataset, ExioBase 3.8.2 (Stadler et al., 2018; Stadler, 2021), a comprehensive source database covering 417 emission categories and 662 material and resource categories occupying over 40 GB of storage. ExioBase 3.8.2 is available in two versions based on different sector classifications. Specifically, the first version is defined on a product-by-product (PxP) basis, covering 200 products, while the second version is defined on an industry-by-industry (IxI) basis, encompassing 163 industries. ExioML treats products and industries as different sector objects, describes them using common features, and creates PxP and IxI into two complementary datasets. Therefore, ExioML preserves the comprehensive information of raw data sources. ExioML provides PxP and IxI data under a similar distribution for self-contained validation. It enables future ML research to evaluate the generalisation and robustness of ML models on two datasets and reduces the effort to seek datasets under the same structures.
@@
Additionally, we open-sourced the well-documented ExioML development toolkit to stimulate interdisciplinary corporations in three contributions. Firstly, by providing the footprint calculation, the toolkit reduces the barrier for new ML researchers without EE domain knowledge. Secondly, data storage and computational demands are substantial due to high-dimensional matrix operations. ExioML development toolkit utilises GPU accelerations and novelty stores multi-dimensional networks in a single edge table, significantly reducing the computing and storage. Thirdly, the toolkit provides a highly flexible interface for querying the factors of interest from ExioBase 3.8.2. MRIO features contain specific Eco-economics meanings, such as various emissions types. Independent features should be carefully selected. Therefore, four essential factors are identified by the EE research community: GHG emissions, population, gross domestic product (GDP), and total primary energy supply (TPES). The decomposition uses the Kaya Identity, a popular structural decomposition method designed specifically for climate change (Peters et al., 2017). This method breaks down carbon emissions into four indicators, which are population ($P$), GDP per capita ($G/P$), energy intensity ($E/G$) and carbon intensity ($F/E$):
@@
4 Technical Validation
GHG emissions are a crucial socio-economic measure for evaluating sectoral sustainability, yet they remain highly uncertain in terms of quantification (Sun & Huang, 2022; Riahi et al., 2007). Emissions are an indirect indicator, unlike value-added, employment, or energy consumption, which can be directly measured. The calculation of sectoral emissions depends on the efficiency of raw materials and production systems, which vary widely across different sectors. Consequently, sectors struggle to disclose their carbon emissions accurately, impacting carbon transparency (Matisoff, 2013). Therefore, our study demonstrates a multimodel sectoral GHG emissions regression task leveraging categorical and numerical features to validate the usability of the proposed ExioML. The model aims to automatically learn the underlying relationships of Eco-economics factors to quantify sectoral GHG emissions.
Regression models could be categorized into shallow and deep learning models. Deep learning in Eco-economics analysis, particularly on tabular data, remains relatively underexplored (Gorishniy et al., 2021). We rigorously evaluated different learning algorithms on tabular learning algorithms, including tree-based models, MLP and Transformer architectures (Vaswani et al., 2017). The distinction lies in their handling of categorical features: shallow models employ Leave-One-Out Text Encoding due to their inability to process categorical data natively (Zhang, 2016), whereas deep learning models learn categorical embeddings directly from raw data. We implemented the deep learning algorithms by PyTorch Tabular package (Joseph, 2021). Algorithms included in GHG emission regression are as follows:
  • K-Nearnest-Neighbour (KNN): This non-parametric method predicts based on the majority vote from its k nearest neighbours (Cover & Hart, 1967).
  • Ridge Regression (Ridge): A linear regression variant employing $\ell_{2}$ regularization to prevent overfitting (Hoerl & Kennard, 1970).
  • Decision Tree (DT): A feature splitting algorithm to tree flowchart structure for decision making (Quinlan, 1986).
  • Random Forest (RF): Ensemble of multiple decision trees for robust prediction (Breiman, 2001).
  • Grandient Boost Decision Tree (DBDT): Ensemble weak trees to improve the error of the previous tree sequentially (Friedman, 2001).
  • Multi-Layer Perceptron (MLP): A deep artificial neural network employing backpropagation with the capability of non-linear function approximation (LeCun et al., 2015).
  • Gated Adaptive Network (GANDALF): GANDALF leverages Gated Feature Learning Units (GFLU) for automated feature selection (Joseph & Raj, 2022).
  • Feature-Tokenizer Transformer (FTT): Novel transformer architecture applies a tokenization mechanism on tabular features (Gorishniy et al., 2021).
Table 2: Impletement detail.
Data Description Setting Description
# Train 141,635 (PxP), 114,678 (IxI) Epoch 30
# Validation 35,409 (PxP), 28,670 (IxI) Batch size 512
# Test 44,262 (PxP), 35,837 (IxI) Shallow Sci-kit Learn
Encoding Shallow models only Deep PyTorch Tabular
Normalisation DL models only Version Python 3.11.7
Scaling Log-transformation GPU Apple M3 Pro 18 GB
PxP IxI
Model MSE Time (s) MSE Time (s)
KNN 1.071 ± 0.010 0.035 ± 0.001 1.151 ± 0.018 0.026 ± 0.001
Ridge 2.265 ± 0.000 0.005 ± 0.002 2.514 ± 0.000 0.004 ± 0.002
DT 0.926 ± 0.051 0.316 ± 0.017 0.848 ± 0.070 0.254 ± 0.019
RF 0.356 ± 0.004 21.521 ± 0.157 0.302 ± 0.004 16.511 ± 0.060
GBDT 0.234 ± 0.006 30.276 ± 0.224 0.219 ± 0.007 32.847 ± 0.388
MLP 0.226 ± 0.007 219.218 ± 0.904 0.250 ± 0.092 205.051 ± 24.309
GANDALF 0.204 ± 0.010 352.756 ± 7.036 0.189 ± 0.007 383.119 ± 3.664
FTT 0.330 ± 0.007 330.578 ± 1.527 0.302 ± 0.023 468.911 ± 7.329
5 Conclusion
Appendix A Competing interests
The authors declare no competing interests.
Appendix B Author Contributions
Y. Guo designed the study and produced the dataset, visualisations, and regression models for technical validation. J. Ma supervised the project. Both authors participated in writing the paper. T. Guan participated in the project design discussion and helped improve the paper draft.
Appendix C Code Availability
Appendix D Data Records & Visualisation
@@
Storing large multi-dimensional Footprint Networks poses a significant challenge. For example, a single factor’s connections for one year amount to 63,792,169, consuming approximately 1 GB of storage. We addressed this challenge by leveraging the characteristics of the global trading network. By representing the network in an edge table, we stored the Multi-dimensional Footprint Networks in a single file, significantly reducing computing and storage requirements. Given that the global trading network is scale-free and edge weights follow a power-law distribution (Foster, 2005), the network is overall sparse, with connections concentrated around dominant nodes. We transformed the graph adjacency matrix into an edge table in tabular format, selecting the top 1,000,000 connections per year. After performing an inner join with the selected factors, around 300,000 edges per year remained. The total storage consumption for the PxP and IxI Footprint Networks is 1.43 GB and 1.34 GB, respectively. The Factor Accounting table requires even less computation and storage, consuming approximately 20 MB.
Appendix E Hyperparamter Tuning

### v2

@@ Abstract
The Environmental Extended Multi-Regional Input-Output analysis is the predominant framework in Ecological Economics for assessing the environmental impact of economic activities. This paper introduces ExioML, the first Machine Learning benchmark dataset designed for sustainability analysis, aimed at lowering barriers and fostering collaboration between Machine Learning and Ecological Economics research. A crucial greenhouse gas emission regression task was conducted to evaluate sectoral sustainability and demonstrate the usability of the dataset. We compared the performance of traditional shallow models with deep learning models, utilizing a diverse Factor Accounting table and incorporating various categorical and numerical features. Our findings reveal that ExioML, with its high usability, enables deep and ensemble models to achieve low mean square errors, establishing a baseline for future Machine Learning research. Through ExioML, we aim to build a foundational dataset supporting various Machine Learning applications and promote climate actions and sustainable investment decisions.
keywords:
Environmental Extended Multi-regional Input-Output, Machine Learning, Sustainability Development, Global Trading Network
1 Background & Summary
@@
The increase in Greenhouse Gas (GHG) emissions due to fossil-fuel-driven economic development has precipitated a global warming crisis. This concern led to establishing the Paris Agreement in 2015, aiming to limit long-term temperature rise to no more than 2 C above pre-industrial levels [1]. Concurrently, the Sustainable Development Goals (SDGs), specifically Goal 13 and Goal 8, were proposed to emphasize taking climate action to reduce emissions while maintaining economic growth. To address such climate-trade dilemma, researchers from various disciplines strive to balance climate action with economic growth and human well-being [2, 3]. Recently, the Machine Learning (ML) technique [4] has emerged as a significant tool for accurate prediction to assist climate change decision-making [5]. Specifically, ML algorithms have been explored in aiding nearly real-time global weather forecasting [6], land monitoring via satellite imagery [7, 8], and the prediction of disturbances in electric grids [9].
The predominant Ecological Economics (EE) research framework, the Environmentally Extended Multiregional Input-Output (EE-MRIO) analysis, effectively models global economic interactions of sectors within a network structure [10, 11, 12, 13, 14]. As EE-MRIO describes the environmental footprint for global economic activities, it has become the fundamental framework for EE research, illustrated in Figure 1. EE-MRIO supports various studies such as Structure Decomposition Analysis (SDA) and Index Decomposition Analysis (IDA) to identify changes by decomposing key drivers [15, 16, 17]; monitoring embodied emissions with resource transfer and sustainability evaluation of supply chain in global trade [18, 19, 20, 21, 22, 23, 12]. Recently, ML algorithms have been applied with EE-MRIO for several applications, such as accurately identifying ecological hotspots and inefficiencies within the global supply chain to optimise logistic paths for decreased carbon emissions while considering cost-effectiveness [24]. ML algorithms are naturally suitable for learning multi-dimensional patterns and are utilised for better sectoral sustainability assessment considering environmental, economic and social impacts [25, 26]. Additionally, the ML algorithms extract the patterns from the original high-dimensional space to latent feature space, and it is leveraged to extract the critical paths of products’ energy, water, and emission footprints from giant trading networks for supporting sustainable investment decisions [27].
@@
However, unlike the open-source culture of the ML community [9], most EE studies haven’t publicised the code and data. This leads to fragmented, independent, and inconsistent research leveraging ML techniques to explore the sustainability of different sectors and regions. Subsequent research cannot fairly compare the model performance, such as accuracy, robustness and generalisation or reproduce the result based on previous research, leading to the slow EE-ML development [28, 29, 30]. Furthermore, no public benchmark EE-ML dataset is available to the best of our knowledge. The EE-MRIO data is a high-quality source for constructing ML-ready datasets. However, the existing dataset faces limitations in terms of accessibility and resolution. Specifically, Eora offers high-quality data but is closed-source and requires high-cost purchasing [31]. Global Trade Analysis Project (GTAP) covers only five reference time steps with non-free access [32]. The World Input-Output Database (WIOD) [33] has been questioned for its low resolution of temporal and spatial scope, whose latest temporal coverage is 2016. Therefore, uniform problem formulations and benchmark datasets need to be urgently developed. In summary, there exists a big gap in the cooperation between Ecological Economics and Machine Learning (EE-ML), which is demonstrated in three aspects: the inaccessibility of the dataset, the challenge of intensive data preprocessing as the domain knowledge being required and the lack of Benchmark ML datasets and models.
To fill these gaps, we proposed ExioML as the first ML-ready benchmark data in EE research. The ExioML is developed on top of the high-quality open-source EE-MRIO dataset ExioBase 3.8.2 with high spatiotemporal resolution [34]. Enabling tracking resource flow between international sectors for global sectoral sustainability assessment. The ExioML developing toolkit is open-sourced, provides Footprint Networks calculation with GPU acceleration, and contributes to the high flexibility of customising interested factors, significantly reducing the barrier for ML research. The usability of the ExioML dataset was validated using multimodality regression to quantify the sectoral GHG emission, achieving a low MSE result. This establishes a critical baseline for accuracy and efficiency in future ML research.
@@ 2 Methods
The architecture of the proposed ExioML dataset is illustrated in Figure 2. It ends the challenges of data limitation and simplifies the EE-MRIO framework for ML researchers in a ready-to-use manner. ExioML offers high spatiotemporal resolution for 49 regions from 1995 to 2022 containing tabular and graphical formats, supporting diverse ML research, such as time-series emissions forecasting [35], factor analysis [36, 37, 17, 38, 39], graph learning [40] and clustering analysis [41, 42], under a uniform framework. The details of ExioML are summarised in Tablue 1.
@@
ExioML is derived from an open-source EE-MRIO dataset, ExioBase 3.8.2 [34, 43], a comprehensive source database covering 417 emission categories and 662 material and resource categories occupying over 40 GB of storage. ExioBase 3.8.2 is available in two versions based on different sector classifications. Specifically, the first version is defined on a product-by-product (PxP) basis, covering 200 products, while the second version is defined on an industry-by-industry (IxI) basis, encompassing 163 industries. ExioML treats products and industries as different sector objects, describes them using common features, and creates PxP and IxI into two complementary datasets. Therefore, ExioML preserves the comprehensive information of raw data sources. ExioML provides PxP and IxI data under a similar distribution for self-contained validation. It enables future ML research to evaluate the generalisation and robustness of ML models on two datasets and reduces the effort to seek datasets under the same structures.
Additionally, we open-sourced the well-documented ExioML development toolkit to stimulate interdisciplinary corporations in three contributions. Firstly, by providing the footprint calculation, the toolkit reduces the barrier for new ML researchers without EE domain knowledge. Secondly, data storage and computational demands are substantial due to high-dimensional matrix operations. ExioML development toolkit utilises GPU accelerations and novelty stores multi-dimensional networks in a single edge table, significantly reducing the computing and storage. Thirdly, the toolkit provides a highly flexible interface for querying the factors of interest from ExioBase 3.8.2. MRIO features contain specific Eco-economics meanings, such as various emissions types. Independent features should be carefully selected. Therefore, four essential factors are identified by the EE research community: GHG emissions, population, gross domestic product (GDP), and total primary energy supply (TPES). The decomposition uses the Kaya Identity, a popular structural decomposition method designed specifically for climate change [16]. This method breaks down carbon emissions into four indicators, which are population ($P$), GDP per capita ($G/P$), energy intensity ($E/G$) and carbon intensity ($F/E$):
@@
3 Data Records
@@
Storing large multi-dimensional Footprint Networks poses a significant challenge. For example, a single factor’s connections for one year amount to 63,792,169, consuming approximately 1 GB of storage. We addressed this challenge by leveraging the characteristics of the global trading network. By representing the network in an edge table, we stored the Multi-dimensional Footprint Networks in a single file, significantly reducing computing and storage requirements. Given that the global trading network is scale-free and edge weights follow a power-law distribution [44], the network is overall sparse, with connections concentrated around dominant nodes. We transformed the graph adjacency matrix into an edge table in tabular format, selecting the top 1,000,000 connections per year. After performing an inner join with the selected factors, around 300,000 edges per year remained. The total storage consumption for the PxP and IxI Footprint Networks is 1.43 GB and 1.34 GB, respectively. The Factor Accounting table requires even less computation and storage, consuming approximately 20 MB.
4 Technical Validation
GHG emissions are a crucial socio-economic measure for evaluating sectoral sustainability, yet they remain highly uncertain in terms of quantification [45, 46]. Emissions are an indirect indicator, unlike value-added, employment, or energy consumption, which can be directly measured. The calculation of sectoral emissions depends on the efficiency of raw materials and production systems, which vary widely across different sectors. Consequently, sectors struggle to disclose their carbon emissions accurately, impacting carbon transparency [47]. Therefore, our study demonstrates a multimodel sectoral GHG emissions regression task leveraging categorical and numerical features to validate the usability of the proposed ExioML. The model aims to automatically learn the underlying relationships of Eco-economics factors to quantify sectoral GHG emissions.
Regression models could be categorized into shallow and deep learning models. Deep learning in Eco-economics analysis, particularly on tabular data, remains relatively underexplored [48]. We rigorously evaluated different learning algorithms on tabular learning algorithms, including tree-based models, MLP and Transformer architectures [49]. The distinction lies in their handling of categorical features: shallow models employ Leave-One-Out Text Encoding due to their inability to process categorical data natively [50], whereas deep learning models learn categorical embeddings directly from raw data. We implemented the deep learning algorithms by PyTorch Tabular package [51]. Algorithms included in GHG emission regression are as follows:
  • K-Nearnest-Neighbour (KNN): This non-parametric method predicts based on the majority vote from its k nearest neighbours [52].
  • Ridge Regression (Ridge): A linear regression variant employing $\ell_{2}$ regularization to prevent overfitting [53].
  • Decision Tree (DT): A feature splitting algorithm to tree flowchart structure for decision making [54].
  • Random Forest (RF): Ensemble of multiple decision trees for robust prediction [55].
  • Grandient Boost Decision Tree (DBDT): Ensemble weak trees to improve the error of the previous tree sequentially [56].
  • Multi-Layer Perceptron (MLP): A deep artificial neural network employing backpropagation with the capability of non-linear function approximation [4].
  • Gated Adaptive Network (GANDALF): GANDALF leverages Gated Feature Learning Units (GFLU) for automated feature selection [57].
  • Feature-Tokenizer Transformer (FTT): Novel transformer architecture applies a tokenization mechanism on tabular features [48].
Table 3: Impletement detail.
Data Description Setting Description
# Train 141,635 (PxP), 114,678 (IxI) Epoch 30
# Validation 35,409 (PxP), 28,670 (IxI) Batch size 512
# Test 44,262 (PxP), 35,837 (IxI) Shallow Sci-kit Learn
Encoding Shallow models only Deep PyTorch Tabular
Normalisation DL models only Version Python 3.11.7
Scaling Log-transformation GPU Apple M3 Pro 18 GB
@@ 4 Technical Validation
PxP IxI
Model MSE Time (s) MSE Time (s)
KNN 1.071 ± 0.010 0.035 ± 0.001 1.151 ± 0.018 0.026 ± 0.001
Ridge 2.265 ± 0.000 0.005 ± 0.002 2.514 ± 0.000 0.004 ± 0.002
DT 0.926 ± 0.051 0.316 ± 0.017 0.848 ± 0.070 0.254 ± 0.019
RF 0.356 ± 0.004 21.521 ± 0.157 0.302 ± 0.004 16.511 ± 0.060
GBDT 0.234 ± 0.006 30.276 ± 0.224 0.219 ± 0.007 32.847 ± 0.388
MLP 0.226 ± 0.007 219.218 ± 0.904 0.250 ± 0.092 205.051 ± 24.309
GANDALF 0.204 ± 0.010 352.756 ± 7.036 0.189 ± 0.007 383.119 ± 3.664
FTT 0.330 ± 0.007 330.578 ± 1.527 0.302 ± 0.023 468.911 ± 7.329
5 Code Availability
Appendix A Competing interests
The authors declare no competing interests.
Appendix B Author Contributions
Y. Guo designed the study and produced the dataset, visualisations, and regression models for technical validation. J. Ma supervised the project. Both authors participated in writing the paper. C. Guan participated in the project design discussion and helped improve the paper draft.